Find an equation for the inverse of the relation y = 6x + 3: Find an equation for the inverse of the relation y = 6x + 3
Answer by Fombitz(1756)

y=6x+3
First step, interchange y for x and vice versa.
highlight(x)=6*highlight(y)+3
Then solve for y.
x=6y+3
6y=x-3
y=(x-3)/6
The inverse of f(x),
f^(-1)(x)=(x-3)/6